An elastic cord is 70 cm long when a weight of 80 N hangs from it but is 84 cm long when a weight of 200 N hangs from it.

What is the "spring" constant k of this elastic cord?

Express your answer to two significant figures and include the appropriate units.
Answer #1

the |_ Μοώ Given data are 2 = 70cm 22 - 84 cm 800 F = goon s 22= 22-ze, = 84-70 = 14cm = Oolumn we know that the hooks law F